OpenStudy (anonymous):

Which of the following equations represents a line that is parallel to y=5x-4 and passes through the point, (3, 4)?

OpenStudy (anonymous):

@AihberKhan

OpenStudy (aihberkhan):

Can you show me the choices?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

y=5x-11 y=5x+19 y=-1/5x-4 y=5x-7

OpenStudy (susanna):

It's y=5x-11. The lines are parallel, so the slope is the same - 5. The line passes through (3,4). Plug those values into the point slope formula: y-y=m(x-x) and you get the answer.

OpenStudy (tkhunny):

We do not need choices. Simple construct it! You have the slope. parallel to y=5x-4 -- so Slope = 5 not -5 You have a point. passes through the point, (3, 4)? Construct it!

OpenStudy (anonymous):

thanks everybody

OpenStudy (susanna):

sure

OpenStudy (tkhunny):

Thanks for what? Did you do it? Let's see what you get.

OpenStudy (anonymous):

lol bro u wit the extras

OpenStudy (tkhunny):

Next time, you show some of YOUR work.

OpenStudy (anonymous):

next time show yo mouf